### v1

# SafeInt: Shielding Large Language Models from Jailbreak Attacks via Safety-Aware Representation Intervention

Abstract
With the widespread real-world deployment of large language models (LLMs), ensuring their behavior complies with safety standards has become crucial. Jailbreak attacks exploit vulnerabilities in LLMs to induce undesirable behavior, posing a significant threat to LLM safety. Previous defenses often fail to achieve both effectiveness and efficiency simultaneously. Defenses from a representation perspective offer new insights, but existing interventions cannot dynamically adjust representations based on the harmfulness of the queries. To address this limitation while ensuring both effectiveness and efficiency, we propose SafeIntervention (SafeInt), a novel defense method that shields LLMs from jailbreak attacks through safety-aware representation intervention. SafeInt is built on our analysis of the representations of jailbreak samples. It adjusts representation distributions of jailbreak samples through intervention to align them with the representations of unsafe samples while minimizing unnecessary perturbations to jailbreak-irrelevant representations. We conduct comprehensive experiments covering six jailbreak attacks, two jailbreak datasets, and two utility benchmarks. Experimental results demonstrate that SafeInt outperforms all baselines in defending LLMs against jailbreak attacks while largely maintaining utility. Additionally, we evaluate SafeInt against adaptive attacks and verify its effectiveness in mitigating real-time attacks.
SafeInt: Shielding Large Language Models from Jailbreak Attacks via Safety-Aware Representation Intervention
Jiaqi Wu1, Chen Chen1, Chunyan Hou2, Xiaojie Yuan1 1Nankai University 2Tianjin University of Technology wujq@mail.nankai.edu.cn, {nkchenchen, yuanxj}@nankai.edu.cn
1 Introduction
Large Language Models (LLMs) OpenAI et al. (2024); Touvron et al. (2023); Grattafiori et al. (2024) have demonstrated remarkable performance across various domains Zhang et al. (2023); Liu et al. (2023); Wang et al. (2024). With their widespread application in real-world scenarios, LLMs face safety challenges Ferrara (2023); Ji et al. (2023). Although efforts Wang et al. (2023a); Rafailov et al. (2024); Zhou et al. (2023) have been made to align LLMs’ behaviors with human values through carefully designed training strategies, recent studies Zou et al. (2023b); Li et al. (2024); Mehrotra et al. (2024); Liu et al. (2024b) reveal that LLMs can still produce undesirable behaviors when subjected to well-crafted jailbreak attacks, such as the biased generation or potentially harmful responses.
Various defense methods have been proposed to address the growing threat of jailbreak attacks. Prompt-based defenses use instructions Phute et al. (2024); Xie et al. (2023); Zhang et al. (2024) or context Zhou et al. (2024a); Wei et al. (2024) to prevent LLMs from generating harmful content. However, prompt-based methods rely on manually crafted secure prompts and possibly lead to excessive self-censorship Varshney et al. (2024), reducing the helpfulness of LLMs for benign queries. Detection-based defenses compute the perplexity of inputs Alon and Kamfonas (2023) or perturb them Cao et al. (2024) to identify jailbreak prompts. Decoding-based defenses Xu et al. (2024); Liu et al. (2024a) reconstruct a safer output probability distribution through contrastive decoding. However, these methods often lack effectiveness or require additional inference overhead. We aim to defend LLMs against jailbreak attacks from a representation perspective, which provides a more controllable and efficient approach. Previous studies Zou et al. (2023a); Rimsky et al. (2024) have shown the effectiveness of intervening representations to steer LLMs’ behaviors, but such interventions cannot dynamically adjust representations based on whether a query is harmful. This limitation makes it challenging to leverage representations for mitigating jailbreak attacks.
In this paper, we analyze the representations of jailbreak samples on four LLMs. Our analysis uses a classifier as a proxy to investigate whether jailbreak representations are distinguishable and whether the representation distributions of different jailbreak methods are consistent. We derive two observations. First, in both intermediate and later layers of LLMs, the representations of jailbreak samples can be distinguished from those of safe or unsafe samples. Second, the consistency of the representation distributions across different jailbreak methods is observed in all LLMs, and it is generally more pronounced in the intermediate layers.
Building on these observations, we propose SafeIntervention (SafeInt), a novel defense method that shields LLMs from jailbreak attacks via safety-aware representation intervention. Since the representations of jailbreak samples differ from those of unsafe samples, LLMs fail to reject jailbreak samples. The core idea of SafeInt is to adjust representation distributions of jailbreak samples through intervention and align them with unsafe representations while minimizing unnecessary perturbations to jailbreak-irrelevant representations. To achieve this, SafeInt adopts a parameterized intervention at an intermediate layer and performs representation alignment in subsequent layers by incorporating classification probabilities and contrastive learning. For jailbreak-irrelevant representations, we perform representation reconstruction to maintain their stability. After training, SafeInt can adaptively identify and intervene in jailbreak-related representations while seamlessly integrating into the LLM inference process.
We conduct a comprehensive evaluation of SafeInt, covering six jailbreak attacks, two jailbreak datasets, and two utility benchmarks. Experimental results show that SafeInt consistently outperforms all baselines in defending against jailbreak attacks. In most cases, SafeInt also maintains the best utility. Additionally, we evaluate SafeInt against adaptive attacks and verify the effectiveness of SafeInt in defending against real-time attacks. In summary, our main contributions are as follows:
  • We observe that the representations of jailbreak samples are distinguishable and that the representation distributions of different jailbreak methods exhibit consistency.
  • We propose SafeInt, a novel defense method that can adaptively identify and intervene in jailbreak-related representations to shield LLMs from jailbreak attacks.
  • Extensive experiments show that SafeInt significantly outperforms all baselines in defending against jailbreak attacks while largely maintaining utility.
2 Preliminaries
2.1 Representation Intervention
Representation intervention is an effective means of steering LLM behavior. For a given decoder-only transformer model with $L$ layers, we denote the internal representation (or residual stream activation) of the last token at layer $l$ as $\mathbf{h}^{(l)}\in\mathbb{R}^{d}$. A typical form of representation intervention is as follows:
$\widetilde{\mathbf{h}}^{(l)}=\mathbf{h}^{(l)}\pm\epsilon\cdot\mathbf{v}.$ (1)
Here, $\widetilde{\mathbf{h}}^{(l)}$ is the intervened representation, $\epsilon\in\mathbb{R}$ represents the intervention strength, and $\mathbf{v}\in\mathbb{R}^{d}$ denotes the intervention direction.
2.2 Analysis of Jailbreak Sample Representations
Recent works have investigated the representation distributions of unsafe and safe samples within LLMs, utilizing their distributional characteristics to enhance safety or facilitate jailbreaks. In this paper, we analyze the representation distributions of three types of samples after introducing jailbreak samples. We construct three training datasets: $\mathcal{D}_{\text{jailbreak}}$, which consists of jailbreak instructions generated only using GCG Zou et al. (2023b) on AdvBench Zou et al. (2023b); $\mathcal{D}_{\text{unsafe}}$, which includes harmful instructions extracted from MaliciousInstruct Huang et al. (2023) and TDC2023 Mazeika et al. (2023); and $\mathcal{D}_{\text{safe}}$, which contains harmless instructions sampled from Alpaca Taori et al. (2023). [^1] For convenience, we abbreviate these datasets as $\mathcal{D}_{\text{j}}$, $\mathcal{D}_{\text{u}}$, and $\mathcal{D}_{\text{s}}$ in the following. More details of the datasets are provided in Appendix A.1. We conduct our analysis on four LLMs: Qwen-7B-Chat Bai et al. (2023), Llama-2-7B-Chat Touvron et al. (2023), Llama-3-8B-Instruct Grattafiori et al. (2024), and Vicuna-7B-v1.5 Chiang et al. (2023). In each layer of the LLM, we train a logistic regression classifier to fit the representations of the three types of samples and report the test accuracy.
Figure 1: Test accuracy of classifiers at different layers of LLMs, with the test set containing only GCG jailbreak samples.
Figure 2: Test accuracy of classifiers on a test set containing jailbreak samples from GCG, AutoDAN, and DeepInception.
Q1: Are the representations of jailbreak, unsafe, and safe samples distinguishable?
In Figure 1, we present the classification accuracy on a test set containing only GCG jailbreak samples. Using the classifier as a proxy for observation, a higher classification accuracy indicates that the representations of the three types of samples are more distinguishable. For all LLMs, the test accuracy remains above 95% starting from the 10th or 11th layer. This indicates that the representations of the three types of samples become distinguishable from the intermediate layers of LLMs onward.
Q2: Are the representation distributions of samples generated by different jailbreak methods consistent?
We reconstruct a test set where jailbreak samples are composed of three methods: GCG, AutoDAN Liu et al. (2024b), and DeepInception Li et al. (2024). We employ the previously trained classifiers for testing and show the results in Figure 2. Since the classifiers are trained solely on GCG jailbreak samples, a high test accuracy reveals that the representations generated by different jailbreak methods exhibit a unified pattern from the classifier’s perspective, indicating consistency.
We observe this consistency across different LLMs. For Qwen, Llama2, and Llama3, the accuracy remains above 90% in most layers. For Vicuna, the accuracy exceeding 90% is primarily observed in the intermediate layers. Although the trend of consistency across layers varies among different LLMs, it is generally more pronounced in the intermediate layers.
Key Insights and Motivation Aligned LLMs can reject unsafe instructions, but due to the distinction in representations between jailbreak samples and unsafe samples, LLMs fail to reject jailbreak instructions, resulting in the generation of harmful content. Based on the distinguishability of jailbreak sample representations and their observed distributional consistency, we aim to intervene in these representations to align their distribution with that of unsafe samples. This alignment enables the LLM to recognize jailbreak samples as unsafe and reject them accordingly.
Figure 3: The schematic of SafeInt. We apply the intervention (illustrated by the blue shield) at a specific layer and perform alignment in the subsequent layers. The distribution of jailbreak sample representations is adjusted to align with that of unsafe samples while minimizing shifts in the representations of safe and unsafe samples. With the original representation distribution, the LLM is successfully jailbroken and generates harmful content. After alignment, the LLM safely rejects the jailbreak instruction.
3 Method
In this section, we describe how SafeInt enhances the safety of LLMs. Figure 3 illustrates the diagram of SafeInt.
3.1 Intervention Parameterization
We aim to find an intervention that, when applied to the representation of the last token in a jailbreak sample, aligns its representation with that of unsafe samples. At the same time, the intervention must not affect the normal responses of safe samples. To this end, we parameterize the intervention. Specifically, we use LoReFT Wu et al. (2024) to parameterize the intervention applied to the last token at a specific layer. LoReFT is an efficient representation fine-tuning method that intervenes on representations within a linear subspace defined by a low-rank projection matrix. Assuming the intervention is applied at layer $\mathcal{I}$, it can be parameterized as follows:
$\widetilde{\mathbf{h}}^{(\mathcal{I})}=\mathbf{h}^{(\mathcal{I})}+\mathbf{U}^{\top}(\mathbf{Wh^{(\mathcal{I})}}+\mathbf{b}-\mathbf{Uh^{(\mathcal{I})}}).$ (2)
The matrix $\mathbf{U}\in\mathbb{R}^{r\times d}$ has orthonormal rows, where $r$ denotes the rank of the subspace, and $\mathbf{W}\in\mathbb{R}^{r\times d}$ and $\mathbf{b}\in\mathbb{R}^{r}$ are the projection matrix and bias vector, respectively.
Then, we define the objectives for learning the intervention. Broadly, our objectives are twofold: a safety objective and a utility objective. The safety objective guarantees the intervention to help the LLM reject jailbreak and harmful instructions. The utility objective ensures that the intervention does not degrade the response quality for harmless instructions, maintaining the LLM’s high availability.
3.2 Representation Alignment
We first use the classifier as a proxy to assess whether the distributions of jailbreak samples and unsafe samples are consistent in the representation space. From the perspective of the classifier, the alignment is achieved when the classification results for jailbreak and unsafe sample representations are consistent. Specifically, for jailbreak samples, we intervene on their representations to maximize the probability of being classified as unsafe. For unsafe sample representations, they should still be classified as unsafe with a high probability.
We denote the sets of original representations of $\mathcal{D}_{\text{j}}$, $\mathcal{D}_{\text{u}}$, and $\mathcal{D}_{\text{s}}$ as $H_{\text{j}}$, $H_{\text{u}}$, and $H_{\text{s}}$, respectively. The sets of intervened representations are denoted as $\widetilde{H}_{\text{j}}$, $\widetilde{H}_{\text{u}}$, and $\widetilde{H}_{\text{s}}$. Let $\mathbb{L}^{a}$ be the set of layers to be aligned, with $\min(\mathbb{L}^{a})>\mathcal{I}$. After applying the intervention at the layer $\mathcal{I}$, the updated representation is propagated to the subsequent layers. At layer $l\in\mathbb{L}^{a}$, we extract the latest representation $\widetilde{\mathbf{h}}^{(l)}$ and compute the following:
$\mathcal{L}_{cls}^{(l)}=$ $-\frac{1}{|\widetilde{H}^{(l)}_{\text{j}}|}\sum_{\widetilde{\mathbf{h}}^{(l)}_{\text{j}}\in\widetilde{H}^{(l)}_{\text{j}}}\log P_{\text{u}}(\widetilde{\mathbf{h}}^{(l)}_{\text{j}})$
$-\frac{1}{|\widetilde{H}^{(l)}_{\text{u}}|}\sum_{\widetilde{\mathbf{h}}^{(l)}_{\text{u}}\in\widetilde{H}^{(l)}_{\text{u}}}\log P_{\text{u}}(\widetilde{\mathbf{h}}^{(l)}_{\text{u}}),$ (3)
where $P_{\text{u}}$ represents the probability that classifier $P$ classifies a representation as unsafe.
Contrastive Learning Although we align the representations of jailbreak samples with unsafe samples by a classifier, the limited training data may prevent the classifier’s decision boundary from accurately capturing the discriminative boundary within the LLM. To enhance the alignment, we use contrastive learning as a complementary task. Contrastive learning is a widely used technique that pulls similar samples closer while pushing different samples apart in the representation space Sohn (2016). For a given representation $\mathbf{q}$, contrastive learning maximizes the similarity between $\mathbf{q}$ and the set of positive samples $K^{+}$ while minimizing the similarity between $\mathbf{q}$ and the set of negative samples $K^{-}$, with the objective formulated as follows:
CT $(\mathbf{q},K^{+},K^{-})=$
$-\log\frac{\exp(\text{sim}(\mathbf{q},\mathbf{k}^{+})/\tau)}{\sum_{\mathbf{k}\in(K^{+},K^{-})}\exp(\text{sim}(\mathbf{q},\mathbf{k})/\tau)},$ (4)
where $\mathbf{k}^{+}\in K^{+}$, $\text{sim}(\cdot,\cdot)$ represents cosine similarity, and the temperature is set to $\tau=0.1$.
Specifically, the intervened representations of jailbreak samples should be as close as possible to those of unsafe samples while being pushed away from their original representations and those of safe samples. Accordingly, for $\widetilde{\mathbf{h}}^{(l)}_{\text{j}}\in\widetilde{H}^{(l)}_{\text{j}}$, the contrastive loss is calculated as:
$\mathcal{L}_{ct}^{(l)}=\text{CT}(\widetilde{\mathbf{h}}^{(l)}_{\text{j}},H^{(l)}_{\text{u}},(H^{(l)}_{\text{j}}\cup H^{(l)}_{\text{s}})).$ (5)
3.3 Representation Reconstruction
To prevent excessive intervention from distorting the LLM’s internal representations, we introduce a reconstruction loss to constrain jailbreak-irrelevant representations from changing. Specifically, we encourage the representations of safe and unsafe samples after intervention to remain close to their original states. This ensures that the intervention primarily affects jailbreak-related representations without causing unnecessary shifts in the model’s overall representation structure. The loss is formulated as follows:
$\mathcal{L}_{recon}=\text{MSE}(H_{\text{s}},\widetilde{H}_{\text{s}})+\text{MSE}(H_{\text{u}},\widetilde{H}_{\text{u}}),$ (6)
where MSE refers to the mean squared error loss.
Considering both the alignment and reconstruction objectives, our final loss is calculated as follows:
$\mathcal{L}_{total}=\alpha\sum_{l\in\mathbb{L}^{a}}(\mathcal{L}_{cls}^{(l)}+\mathcal{L}_{ct}^{(l)})+\beta\mathcal{L}_{recon}.$ (7)
Through the two hyperparameters $\alpha$ and $\beta$, we achieve a balance between effective alignment and model stability.
4 Experiments
4.1 Experimental Setup
Models and Datasets We evaluate SafeInt on two open-source LLMs: Llama2-7b-chat and Vicuna-7b-v1.5. To assess the effectiveness of our defense against jailbreak attacks, we randomly sample 50 instructions from AdvBench Zou et al. (2023b) as the test set, ensuring no overlap with the training set $\mathcal{D}_{\text{j}}$. Additionally, to demonstrate that SafeInt is data-agnostic, we construct an out-of-distribution test set consisting of 50 instructions randomly sampled from JailbreakBench Chao et al. (2024a). Beyond defense, it is also essential to consider the impact on generation quality. Following Xu et al. (2024), we use MT-Bench Zheng et al. (2023) and Just-Eval Lin et al. (2023) to assess the utility of intervened LLMs.
Jailbreak Attacks Multiple representative jailbreak attacks are employed in our evaluation. These include optimization-based attacks: GCG and AutoDAN, LLM-generated attacks: PAIR Chao et al. (2024b) and TAP Mehrotra et al. (2024), and scenario-based attacks: DeepInception. We also consider multilingual mismatch generalization attacks (MG) Yong et al. (2024), where each instruction in the test set is translated into one of six non-English languages to perform the attacks. Detailed setup can be found in Appendix A.2.
Baselines We compare SafeInt with six state-of-the-art defense approaches: PPL Alon and Kamfonas (2023), Paraphrase Jain et al. (2023), Self-Examination Phute et al. (2024), ICD Wei et al. (2024), Self-Reminder Xie et al. (2023), and SafeDecoding Xu et al. (2024).
Evaluation Metrics We use two types of Attack Success Rate (ASR) to evaluate defense effectiveness: ASR-keyword, which matches predefined refusal keywords, and ASR-GPT, which leverages GPT-4o-mini to assess whether the LLM generates harmful content relevant to the malicious instruction. Lower ASR values indicate better defense performance. For MT-Bench and Just-Eval, we adopt GPT-based scoring, where Just-Eval evaluates five aspects: helpfulness, clarity, factuality, depth, and engagement.
Implementation Details Previous classification results indicate that in the intermediate layers, the representations of various jailbreak samples are relatively consistent, and jailbreak samples can be effectively distinguished from both unsafe and safe samples. To avoid time-consuming searches, we directly select layer $\mathcal{I}=12$ as the intervention layer. Since Vicuna exhibits weaker security, deeper alignment is required, so we set its alignment layers from layer $15$ to the final layer. For Llama2, which has undergone safety reinforcement, aligning only the final layer is sufficient.
Model Defense Benchmark $\downarrow$ Jailbreak Attacks $\downarrow$
AdvBench GCG AutoDAN DeepInception PAIR TAP MG
Vicuna No Defense 8% (4%) 90% (92%) 82% (88%) 64% (100%) 54% (60%) 84% (80%) 30% (66%)
PPL 8% (4%) 26% (30%) 72% (68%) 64% (100%) 52% (58%) 84% (82%) 28% (62%)
Paraphrase 6% (6%) 18% (20%) 34% (52%) 38% (96%) 36% (38%) 42% (52%) 10% (32%)
Self-Examination 2% (0%) 12% (16%) 18% (22%) 34% (74%) 8% (14%) 34% (30%) 6% (34%)
ICD 0% (0%) 14% (14%) 40% (36%) 64% (96%) 24% (34%) 44% (44%) 6% (34%)
Self-Reminder 0% (0%) 4% (6%) 8% (6%) 46% (100%) 26% (32%) 38% (40%) 16% (50%)
SafeDecoding 0% (0%) 2% (2%) 10% (4%) 0% (0%) 4% (6%) 12% (12%) 12% (40%)
SafeInt (Ours) 0% (0%) 0% (0%) 2% (2%) 0% (0%) 2% (6%) 8% (10%) 4% (8%)
Llama2 No Defense 0% (0%) 30% (32%) 34% (44%) 0% (0%) 2% (10%) 10% (10%) 0% (6%)
PPL 0% (0%) 0% (2%) 2% (8%) 0% (0%) 2% (8%) 10% (10%) 0% (4%)
Paraphrase 0% (10%) 0% (22%) 6% (26%) 0% (0%) 2% (30%) 2% (30%) 0% (16%)
Self-Examination 0% (0%) 0% (4%) 2% (6%) 0% (0%) 2% (4%) 2% (4%) 0% (0%)
ICD 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (0%)
Self-Reminder 0% (0%) 0% (2%) 0% (0%) 0% (0%) 2% (4%) 0% (2%) 0% (6%)
SafeDecoding 0% (0%) 0% (2%) 0% (4%) 0% (0%) 0% (6%) 0% (0%) 0% (0%)
SafeInt (Ours) 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (4%) 0% (0%) 0% (0%)
Table 1: ASR-GPT (outer) and ASR-keyword (in parentheses) for different defense methods on AdvBench. The best results are in bold. SafeInt outperforms all baselines across various attacks.
Model Defense Benchmark $\downarrow$ Jailbreak Attacks $\downarrow$
JailbreakBench GCG AutoDAN DeepInception PAIR TAP MG
Vicuna No Defense 6% (10%) 74% (96%) 76% (98%) 54% (100%) 42% (46%) 66% (68%) 30% (76%)
PPL 6% (10%) 20% (30%) 48% (62%) 46% (100%) 38% (48%) 66% (68%) 26% (66%)
Paraphrase 6% (20%) 18% (40%) 22% (60%) 28% (98%) 16% (36%) 32% (42%) 10% (40%)
Self-Examination 0% (4%) 8% (28%) 26% (48%) 28% (74%) 10% (14%) 28% (30%) 8% (50%)
ICD 0% (0%) 8% (14%) 42% (42%) 54% (94%) 16% (28%) 32% (42%) 22% (52%)
Self-Reminder 0% (2%) 4% (4%) 4% (6%) 36% (100%) 14% (20%) 26% (30%) 30% (62%)
SafeDecoding 0% (0%) 0% (0%) 18% (18%) 0% (0%) 10% (14%) 10% (12%) 14% (36%)
SafeInt (Ours) 0% (0%) 0% (0%) 4% (6%) 0% (0%) 2% (12%) 4% (12%) 8% (24%)
Llama2 No Defense 0% (0%) 24% (34%) 30% (42%) 2% (2%) 0% (6%) 6% (6%) 0% (4%)
PPL 0% (0%) 2% (2%) 6% (6%) 2% (2%) 0% (4%) 6% (6%) 0% (2%)
Paraphrase 0% (6%) 0% (28%) 0% (22%) 2% (2%) 0% (24%) 4% (28%) 0% (10%)
Self-Examination 0% (0%) 2% (2%) 6% (6%) 0% (0%) 0% (4%) 2% (2%) 0% (2%)
ICD 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (0%)
Self-Reminder 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (10%)
SafeDecoding 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (2%) 0% (6%) 0% (0%)
SafeInt (Ours) 0% (0%) 0% (0%) 0% (0%) 0% (0%) 0% (6%) 0% (0%) 0% (0%)
Table 2: ASR-GPT (outer) and ASR-keyword (in parentheses) on JailbreakBench. The best results are in bold. SafeInt consistently achieves the best performance.
Model Defense MT-Bench $\uparrow$ Just-Eval $\uparrow$
Helpfulness Clear Factual Deep Engaging Average
Vicuna No Defense 5.21 4.44 4.66 4.38 3.60 3.49 4.11
Self-Examination 5.03 4.40 4.65 4.34 3.56 3.47 4.08
ICD 4.86 4.34 4.61 4.34 3.40 3.32 4.00
SafeDecoding 4.84 3.92 4.45 4.19 3.24 3.25 3.81
SafeInt (Ours) 5.09 4.40 4.64 4.35 3.49 3.41 4.06
Llama2 No Defense 5.80 4.65 4.78 4.50 4.19 3.90 4.40
Self-Examination 1.61 3.21 3.67 3.47 2.92 2.68 3.19
ICD 2.91 3.44 4.08 3.96 3.25 3.24 3.59
SafeDecoding 5.68 4.53 4.73 4.42 4.05 3.83 4.31
SafeInt (Ours) 5.82 4.62 4.76 4.47 4.13 3.89 4.37
Table 3: Utility evaluation scores of SafeInt and baselines. The highest and second-highest scores obtained by defense methods are in bold and underlined, respectively. SafeInt maintains the best utility in most cases.
4.2 Main Results
Table 1 presents the ASR results of SafeInt and various baselines on AdvBench. For both Vicuna and Llama2, SafeInt achieves the best performance, reducing ASR to the lowest level among all defense methods under different attacks. Although our training process only utilizes jailbreak samples constructed with GCG, SafeInt effectively defends against other attack strategies, such as PAIR and TAP, which generate adversarial prompts using the LLM. This highlights the generalization capability of our defense, validating our previous observation. Moreover, even against MG attacks, SafeInt significantly lowers ASR, showing that it can generalize to different languages.
Table 2 reports results on another out-of-distribution test set, JailbreakBench. SafeInt continues to outperform all baselines across different models and attack strategies. This demonstrates its robustness to unseen data.
While delivering strong defense performance, SafeInt largely preserves the utility of LLMs. As shown in Table 3, SafeInt achieves almost identical scores to the non-defended model in Llama2, whereas ICD and Self-Examination severely degrade utility. For Vicuna, SafeInt results in only a 2% decrease in MT-Bench and a 1% decrease in Just-Eval compared to the non-defended model. In contrast, SafeDecoding leads to 7% drops in both benchmarks.
Due to the efficiency of LoReFT, SafeInt can be seamlessly integrated into LLM inference, similar to LoRA Hu et al. (2021). Unlike SafeDecoding, which requires an additional expert model for contrastive decoding, SafeInt introduces virtually no extra computational overhead and is therefore more efficient.
Jailbreak Attacks AdvBench JailbreakBench
Adaptive-GCG 0% (0%) 0% (6%)
Adaptive-AutoDAN 0% (0%) 6% (8%)
Table 4: Experimental results of defending against adaptive attacks on Vicuna, with evaluation metrics ASR-GPT and ASR-keyword (in parentheses). ’Adaptive-GCG’ and ’Adaptive-AutoDAN’ refer to GCG and AutoDAN attacks that are optimized in real-time based on the LLM deployed with SafeInt.
4.3 Adaptive Attack
We also consider a scenario where the attacker knows SafeInt and has access to the LLM deployed with it. This means the attacker can dynamically adjust their attack strategies based on the latest defended LLM to bypass the defense mechanisms more effectively. Table 4 presents the experimental results in this scenario, showing that SafeInt maintains strong defensive performance. After SafeInt is deployed, even if GCG and AutoDAN optimize their adversarial prompts in real time, it is difficult to generate threatening attacks.
5 Analyses
5.1 Ablation Studies
We conduct ablation studies on the introduced contrastive loss and reconstruction loss to verify their effectiveness. As shown in Table 5, removing contrastive loss increases ASR, indicating its crucial role in enhancing defense performance. Incorporating contrastive loss leads to a decrease in MT-Bench scores, which may be attributed to its impact on the overall representation structure when pulling together or pushing apart local representations. When the reconstruction loss is omitted, representations are more susceptible to excessive intervention, resulting in both defense failure and a significant decline in response quality.
Methods Jailbreak Attacks $\downarrow$ MT-Bench $\uparrow$
GCG AutoDAN PAIR
No Defense 90% 82% 54% 5.21
SafeInt 0% 2% 2% 5.09
w/o $\mathcal{L}_{ct}$ 2% 8% 6% 5.26
w/o $\mathcal{L}_{recon}$ 2% 12% 8% 4.09
Table 5: Ablation results of our method on AdvBench and MT-Bench, using ASR-GPT as the metric for Jailbreak Attacks. ’w/o $\mathcal{L}_{ct}$’ and ’w/o $\mathcal{L}_{recon}$’ denote the removal of contrastive loss and reconstruction loss, respectively.
(a) Intervention Layer $\mathcal{I}$
(b) Alignment Layer Range $\mathbb{L}^{a}$
Figure 4: Analysis of the intervention layer and alignment layer range.
5.2 Hyperparameter Analysis
Intervention Layer Choice To understand how the choice of the intervention layer impacts our defense effectiveness, we conduct an analysis. Figure 4(4(a)) displays the ASR-keyword when the intervention layer is set between layers 10 and 20. We observe that intervening in the intermediate layers generally yields better results than intervening in the later layers, which may suggest that these intermediate layers play a more crucial role in jailbreak mechanisms. Notably, when the intervention is applied at layer 13, ASR reaches its lowest point. This finding aligns with our observation in Figure 2, where Vicuna exhibits the highest jailbreak representation consistency at layer 13.
Alignment Layer Range We fix the endpoint of the alignment layer range at the final layer and modify the starting point to control its span. In Figure 4(4(b)), we illustrate the results when setting the starting point between layers 15 and 25. We observe that as the starting point shifts to later layers, the defense effectiveness weakens. This may be attributed to the reduced number of aligned layers being insufficient to correct the attack. Overall, while adjusting the alignment layer range impacts defense performance, the effect is not drastic, indicating that our method exhibits a certain degree of robustness to this hyperparameter.
6 Related Work
Jailbreak Attacks Jailbreak attacks aim to bypass alignment or safeguards, forcing LLMs to generate inappropriate content. Early jailbreak attacks Wei et al. (2023); Yong et al. (2024); Yuan et al. (2024) rely on manually crafted adversarial prompts, which primarily exploit objective competition and mismatched generalization to achieve jailbreaks. Subsequent optimization-based attacks Zou et al. (2023b); Liu et al. (2024b); Paulus et al. (2024) introduce automated adversarial prompt optimization by leveraging the internal states of LLMs, significantly improving both the success rate and efficiency of jailbreaks. Recent jailbreak attacks Chao et al. (2024b); Mehrotra et al. (2024); Ding et al. (2024) iteratively rewrite and refine adversarial prompts using one or multiple LLMs, further exposing security vulnerabilities in LLMs.
Jailbreak Defenses To address the challenges posed by jailbreak attacks, numerous defense methods have been proposed Robey et al. (2024); Kumar et al. (2025). Detection-based approaches identify adversarial prompts by computing perplexity Alon and Kamfonas (2023) or randomly deleting parts of the input Cao et al. (2024). Some methods prompt the LLM to perform self-checking through instructions Phute et al. (2024); Xie et al. (2023); Zhang et al. (2024) or context Zhou et al. (2024a). Decoding-based defenses Xu et al. (2024); Liu et al. (2024a) focus on analyzing decoding probabilities under different conditions and formulating decoding strategies to ensure safer outputs. Additionally, certain approaches Zhao et al. (2024); Ouyang et al. (2025) edit specific model parameters to make LLMs forget harmful knowledge. A more controllable and efficient class of defenses Li et al. (2025); Shen et al. (2025) involves manipulating representations to mitigate jailbreak attacks without modifying model parameters or adding decoding overhead.
Representation Engineering for Safety Many studies have employed representation engineering techniques Zou et al. (2023a) to investigate or enhance the safety of LLMs. Zhou et al. (2024b) and Arditi et al. (2024) analyze the mechanisms of jailbreak and refusal from a representation perspective, respectively. Li et al. (2025) improve the robustness of LLMs by strengthening the safety patterns they recognize. Zheng et al. (2024) introduce a learnable safety prompt that aims to increase the separation between harmful and harmless query representations along the refusal direction. Shen et al. (2025) add a difference vector to query representations to guide the LLM toward rejecting malicious instructions, while Gao et al. (2024) mitigate jailbreak attacks by constraining activations within a safe boundary. A major drawback of these two approaches is that their interventions cannot be automatically optimized. This means that when the intervention is applied to all query representations, the choice of intervention strength becomes highly sensitive. In contrast, our method adopts a parameterized intervention, which adaptively identifies and adjusts jailbreak-related representations regardless of manually tuning the intervention strength.
7 Conclusion
This paper first analyzes the representations of jailbreak samples on four LLMs and makes key observations. Building on these observations, we propose SafeIntervention (SafeInt), a novel method that defends LLMs against jailbreak attacks via representation intervention. SafeInt can adaptively identify and intervene in jailbreak-related representations while seamlessly integrating into the LLM inference process. Comprehensive experimental results show that our proposed SafeInt outperforms all baselines in defending against jailbreak attacks. In most cases, SafeInt also achieves the best utility.
Limitations
We discuss the limitations of our work. We make a preliminary observation that SafeInt can transfer to different but homologous LLMs. We speculate that these homologous LLMs may share similar jailbreak representation structures. However, we have not conducted an in-depth exploration of the transferability of SafeInt.
Appendix A Appendix
A.1 Dataset Details
A.1.1 Training Data
To construct $\mathcal{D}^{(\text{train})}_{\text{jailbreak}}$, we use GCG to generate jailbreak instructions from 128 randomly sampled instructions from AdvBench Zou et al. (2023b).
To construct $\mathcal{D}^{(\text{train})}_{\text{unsafe}}$, we sample 128 harmful instructions from MaliciousInstruct Huang et al. (2023) and TDC2023 Mazeika et al. (2023).
To construct $\mathcal{D}^{(\text{train})}_{\text{safe}}$, we sample 128 harmless instructions from Alpaca Taori et al. (2023).
A.1.2 Test Data in Q1
To construct $\mathcal{D}^{(\text{test})}_{\text{jailbreak}}$, we first resample 150 instructions from AdvBench. We then use GCG to generate jailbreak instructions from these 150 instructions.
To construct $\mathcal{D}^{(\text{test})}_{\text{unsafe}}$, we sample 150 harmful instructions from Do Not Answer Wang et al. (2023b), MaliciousInstruct, and TDC2023.222Due to the insufficient data volume of the original two datasets, we introduce Do Not Answer.
To construct $\mathcal{D}^{(\text{test})}_{\text{safe}}$, we sample 150 harmless instructions from Alpaca.
Note that $\mathcal{D}^{(\text{test})}_{\text{jailbreak}}$, $\mathcal{D}^{(\text{test})}_{\text{unsafe}}$, and $\mathcal{D}^{(\text{test})}_{\text{safe}}$ do not overlap with $\mathcal{D}^{(\text{train})}_{\text{jailbreak}}$, $\mathcal{D}^{(\text{train})}_{\text{unsafe}}$, and $\mathcal{D}^{(\text{train})}_{\text{safe}}$.
A.1.3 Test Data in Q2
To construct $\mathcal{D}^{({\scriptscriptstyle\text{OOD}})}_{\text{jailbreak}}$, we first sample 50 instructions from AdvBench. We then use three jailbreak methods—GCG, AutoDAN, and DeepInception—to generate jailbreak instructions from these 50 samples, resulting in a total of 150 jailbreak instructions.
For unsafe and safe samples, we continue to use $\mathcal{D}^{(\text{test})}_{\text{unsafe}}$ and $\mathcal{D}^{(\text{test})}_{\text{safe}}$.
Note that $\mathcal{D}^{({\scriptscriptstyle\text{OOD}})}_{\text{jailbreak}}$ does not overlap with $\mathcal{D}^{(\text{test})}_{\text{jailbreak}}$ or $\mathcal{D}^{(\text{train})}_{\text{jailbreak}}$.
A.2 Attack Setup
For GCG, AutoDAN, DeepInception, PAIR, and TAP, we follow the settings from their original papers. For multilingual mismatch generalization attacks, we translate each instruction in the test set into one of six non-English languages—Chinese, Italian, Vietnamese, Arabic, Korean, or Thai—to carry out the attacks.
A.3 Details of ASR-keyword
Table 6 lists the refusal keywords used to evaluate ASR-keyword. We adopt the settings from Xu et al. (2024) and further refine them by adding "I am not programmed to" and "As a responsible". Additionally, we exclude "I’m an" and "I’m just" to prevent false positives.
I’m sorry
I am sorry
I apologize
As an AI
As a responsible
As a language model
As an assistant
I cannot
I can’t
I am unable to
I’m unable to
I’m just an AI
I do not
It is not</s>
It is not appropriate
I am not allowed to
I am not programmed to
I am an AI language model
Table 6: Refusal Keywords